The reaction of calcium oxide with the pollutant, sulfur trioxide, to produce calcium sulfate has been proposed as one way to remove the SO3 from the smoke resulting from burning high sulfur coal (as in a power plant). Using the values from appendix C in the handout or the text calculate the standard free energy change (G0) for the reaction at 298K.
1
Expert's answer
2020-03-26T03:02:07-0400

The reaction equation is the following:

CaO + SO3 "\rightarrow" CaSO4

According to Hess's law, the free energy change of the reaction can be calculated using the free energies of formation of the products and reagents. I have found the following values of the standard free energies of formation of the components in the internet: (https://www.chem.wisc.edu/deptfiles/genchem/netorial/modules/thermodynamics/table.htm):

"\Delta G_f^o", at 298 K, kJ/mol

CaO -604.03

SO3 -371.06

CaSO4 -1321.79

From these values, the standard free energy change of the reaction is:

"\Delta G^o = \Delta G_f^o(CaSO_4) - \Delta G_f^o(CaO) - \Delta G_f^o(SO_3)"

"\Delta G^o = -1321.79 + 604.03 + 371.06 = -346.7 \text{ kJ/mol}"

Answer: "\Delta G^o =" -346.7 kJ/mol